# Option Pricing with Time-Varying Volatility Risk Aversion

### v3

@@ 5.1 Data
Our empirical analysis is based on daily returns for the S&P 500 index, the CBOE VIX, and the SPX option prices based on the S&P 500 index. Our sample period spans 32 years from January 2nd, 1990, to December 31, 2021, with 8,064 trading days. Returns are defined from cum-dividend logarithmic transformed closing prices, which were downloaded from the Wharton Research Data Services (WRDS). Daily VIX prices were obtained from the Chicago Board Options Exchange (CBOE) website. The SPX option prices were obtained from two sources. Prices for the first six years (1990–1995) are the so-called Optsum data, which were purchased from CBOE website. Option prices for the remaining 26 years are OptionMetrics data downloaded from the WRDS database.
Option prices were preprocessed following Christoffersen et al., (2014) and Bakshi et al., (1997). Specifically, we include out-of-the-money put and call option prices with positive trading volume and maturities between two weeks and six months. Put option prices are converted to call options using the put-call parity. Very deep out-of-the-money options with deltas larger than 0.85 or less than 0.15 are discarded. Much of the existing literature uses weekly data, and a panel of option prices is typically sampled on Wednesdays, because liquidity tends to be highest on Wednesdays. However, in our analysis, we use daily option prices, because a daily score, $s_{t}$, is needed to update $\eta_{t}$. The number of available options has grown rapidly over the sample period, both in terms of available maturities and the range of moneyness at each maturity. [^12] The number of available option prices has increased almost 50-fold over our sample period, initially from about 38 daily option prices to well over 1,700. From the pool of available options, we select up to six options per trading day. The inclusion criteria are as follows: for each maturity, we determine the most liquid option on day $t$, as measured by daily trading volume. We sort these options by maturity in ascending order and index these by $j=1,\ldots,\tilde{N}_{t}$, where $\tilde{N}_{t}$ is the number of distinct maturities on day $t$. From this set, we include all options if $\tilde{N}_{t}\leq 6$; the first six options if $7\leq\tilde{N}_{t}\leq 10$; options $\{1,3,5,7,9,11\}$ if $11\leq\tilde{N}_{t}\leq 15$; and options $\{1,4,7,10,13,16\}$ if $16\leq$ $\tilde{N}_{t}\leq 20$; and so forth. This resulting set of options will be representative for the range of available maturities, and we have $N_{t}=6$ on most days. The total number of option prices in our sample period is 37,152.
@@
Note: Summary statistics for close-to-close S&P 500 index returns, VIX, and option prices from January 1990 to December 2021. We report the sample mean (Mean), standard deviation (Std), skewness (Skew), kurtosis (Kurt), number of observations (Obs), for returns and the VIX. Option prices are based on closing prices of out-of-the-money call and put options. We report the average Black-Scholes implied volatility (IV), average price, and the number of option prices for different ranges of Moneyness, days to maturity (DTM), and VIX levels. Data sources: S&P 500 returns from WRDS, VIX from CBOE’s website. Option prices from Optsum data (1990–1995) and OptionMetrics (1996–2021).
Table 2 contains descriptive statistics for our S&P 500 returns and the VIX in Panel A, and option prices in Panel B. As expected, the sample average of the VIX (19.48%) is larger than the standard deviation for annualized returns (18.11%). This difference reflects the (average) negative volatility risk premium. The S&P 500 returns exhibit slight negative skewness and a high degree of kurtosis, whereas the VIX has positive skewness and slightly lower kurtosis. We also report summary statistics for option prices partitioned by moneyness, maturity, and the contemporaneous level of VIX. Options with deltas below 0.5 are out-of-the-money call options, and options with deltas above 0.5 are based on out-of-the-money put options. Deep out-of-the-money put options (i.e., deltas greater than 0.7) are expensive relative to out-of-the-money call options, reflecting the well-known volatility smirk. The implied volatility has a relatively flat term structure with respect to time to maturity. Unsurprisingly, the implied volatility increases with the VIX level, as shown at the bottom of Table 2, where option prices are partitioned by the contemporaneous level of the VIX.
@@ 5.2 Parameter estimation
We estimate the model with both constant and time-varying variance risk aversion, CHNG and DHNG, respectively. [^13] The corresponding results for the Auxiliary structure are presented in the Appendix B Both specifications are estimated using two types of derivative prices, VIX data and the panel of option prices. Parameters are estimated by maximizing the joint log-likelihood function for the full sample period from January 1990 to December 2021. Each column reports the parameter estimates for the specification listed in the first row of Table 3. The type of derivatives, VIX or option prices, used in the estimation is indicated with [VIX] and [Opt], respectively. Robust standard errors are give in parentheses below the estimates. [^14] Following Christoffersen et al., (2013), we impose $\omega$ = 0 in estimation when the non-negativity constraint, which ensures positive variances, is binding. We also report the implied persistence of volatility under both $\mathbb{P}$ and $\mathbb{Q}$. These are given by $\pi^{\mathbb{P}}=\beta+\alpha\gamma^{2}$ and $\pi^{\mathbb{Q}}=\mathbb{E}^{\mathbb{Q}}[\beta_{t}^{*}+\alpha_{t}^{*}\gamma_{t-1}^{*2}]$, respectively, where the latter simplifies to $\pi^{\mathbb{Q}}=\beta^{*}+\alpha^{*}\gamma^{*2}$ for the CHNG model with constant parameters. We also report the different terms of the log-likelihood (for returns and different sets of derivative prices). Some of these likelihood terms can be interpreted as pseudo out-of-sample log-likelihood values, indicated in italic. For instance, the CHNG[VIX] model is estimated using returns and the VIX, but the estimated model also yields model-based option prices that can be compared with actual option prices. The reported log-likelihood for option prices is evaluated with the resulting option pricing errors that are implicitly used to obtain estimates of $\rho$ and $\sigma_{e}$ for option prices. Similarly, we evaluate the log-likelihood for VIX pricing errors for the specifications estimated with option prices. In this case, we compute the implied estimate of $\sigma_{e}$ for VIX prices. The largest log-likelihood within each row is highlighted in bold.
@@
There are several interesting observations to be made from Table 3. First, the volatility process is found to be highly persistent across all specifications, and the persistence is larger under $\mathbb{Q}$ than under $\mathbb{P}$, which is consistent with the existing literature. Second, the estimate of the equity risk premium, $\lambda$, is positive and significant in all specification. Third, the VRR, $\eta_{t}$, is typically larger than one and is significantly larger than one in both CHNG specifications. Similarly, for the DHNG model the expected value of $\eta_{t}$ (denoted $\mathbb{E}\eta$) is inferred from the AR(1) model for $\log\eta_{t}$, and it is also significantly larger than one in both DHNG specifications. This implies that the risk-neutral volatility, $h^{\ast}$, is (on average) larger than the physical volatility, $h$. Fourth, the parameter that defines the leverage effects, $\gamma$, is also found to be significant across all specification. Fifth, for the DHNG specification we note that, $\eta_{t}$, is highly persistent, as $\varphi$ is estimated to be close to one. In fact, it is estimated to be more persistent than $h_{t}$ and $h_{t}^{\ast}$ across all DHNG specifications. Sixth, the coefficient, $\sigma$, which measures the impact of $s_{t-1}$ on $\eta_{t}$, is estimated to be positive and significant. [^15] In principle, $\sigma$ could be negative, but $\operatorname{var}(\sigma s_{t})=\sigma^{2}$ is non-negative. The implied unconditional variance for $\eta_{t}$ is similar for the two DHNG specifications, $\mathrm{var}(\eta_{t})=0.29$ when estimated with VIX and $\mathrm{var}(\eta_{t})=0.23$ when estimated with option prices. [^16] The variance for $\eta_{t}$ is computed from AR(1) parameters for $\log\eta_{t}$ and the assumptions imply that the unconditional distribution for $\eta_{t}$ is log-normal. Seventh, all specifications yield similar likelihood values for the returns. This is to be expected because they all rely on the same Heston-Nandi GARCH model for returns under physical measure. Eighth, the key difference between the CHNG and DHNG models is the enhanced flexibility in DHNG’s pricing kernel. This generalization leads to large improvements in the likelihood for derivative prices. The reason is that the adaptive pricing kernel with time-varying variance risk aversion yields model-implied derivative prices that are much closer to observed prices, and the substantial reduction in the pricing errors translates to much higher values of the log-likelihood for derivative prices.
Ninth, estimating the models with VIX or option prices results in some differences across the estimated parameters. This is to be expected, as the vector of option prices contains more information about the distribution of future returns than the VIX. For instance, the leverage parameter, $\gamma$, is estimated to be larger with option prices than with the VIX.
@@
Tenth, the correlation between option pricing errors is estimated to be positive and significant for both models. The estimate has the staggering large value of $82.2\%$ for the CHNG model and a more moderate value of $8.7\%$ for DHNG model. Ignoring this correlation (assuming it to be zero) would greatly underestimate the condition variance of $\nabla_{t}$, which makes the variance of $s_{t}$ larger than one.
@@
Figure 5: This figure presents time series of the estimated variance risk ratio, $\eta_{t}$. The blue line is based on the VIX and the red line on option prices. The two time series are very similar, albeit the $\eta_{t}$ based on the VIX tend to be slightly larger than that based on option prices.
@@
Figure 5 presents the estimated time series of the VRR, $\eta_{t}=h_{t+1}^{\ast}/h_{t+1}$, based on the VIX (blue line) and option prices (red line). The two have a high degree of comovement and, as we discussed earlier, the unconditional variance of the two series is very similar. Importantly, the VRR occasionally falls below one, as is seen during the years 1993–1995, 2004–2007, and 2017. A value below one ($\eta_{t}<1$) indicates that investors have an increased appetite for variance risk, whereas a large value of $\eta$ implies that investors demand additional compensation for variance risk. The latter was particularly the case during financial crises, such as the immediate aftermath of the Lehman collapse and the recent COVID-19 pandemic. The observed conditional variance was unusually high during these these episodes, and the large value of $\eta_{t}$ implies that $h_{t}^{\ast}$ increased to far higher levels and was, briefly, more than three times larger than $h_{t}$.
@@
Note: This table reports the in-sample VIX and option pricing performance for each model in Table 3. We evaluate the model’s VIX and option pricing ability through the root mean square error for $\log{\rm VIX}$ and log implied volatility ($\log{\rm IV}$), respectively, see (21) and (22). Results for option pricing are presented for different ranges of moneyness, DTM, and the VIX.
@@ 5.3 In-Sample VIX and Option Pricing Performance
We evaluate the performance of the models in terms of their ability to accurately price the VIX and options. We report the root mean squared errors (RMSE) for the logarithm of VIX and the logarithmically transformed implied volatilities. The RMSE multiplied by 100 can be interpreted as the (absolute) relative pricing errors in percent. This loss function is coherent with our log-likelihood function, where the term that involves derivative prices is (4). [^17] As a robustness check, we also computed the RMSE for the level of VIX and implied volatilities. These results are reported in Appendix B. A similar loss function was adopted in Ornthanalai, (2014), who used the relative implied volatility to prevent days with high volatility from being given a disproportionately large weight in the comparisons. For each of the specifications listed in Table 3, we report their (in-sample) RMSEs in Table 4.
@@
where $\mathrm{VIX}_{t}^{m}$ is the model-based quantity and $\mathrm{VIX}_{t}$ is the observed market VIX. Panel A of Table 4 shows that a time-varying VRR, $\eta_{t}$, results in substantially smaller average pricing errors, relative to a constant $\eta$, which is a characteristic of CHNG. The improved VIX pricing is substantial and impressive. The most accurate VIX pricing is achieved with by the DHNG model, when it is estimated with VIX data. This reduces the RMSE by a factor of four relative to both CHNG specifications. Even the DHNG model that is estimated solely with option prices, has a much smaller RMSE for VIX prices than the CHNG specification that uses VIX prices as part of the objective in the estimation. This is impressive, because the DHNG model estimated with option prices is based on an objective that simultaneously seeks to price options with maturities ranging from one to six months. This entails a trade-off across maturities, unlike models estimated with solely with VIX, that specifically targets the one-month maturity.
@@
Figure 6: Daily model-based (red lines) and market-based (red lines) for the VIX and average implied volatility. The CHNG and DHNG models are estimated with VIX data in the upper panels and option prices in the lower panels.
@@
Note: Each model is estimated once, using data from the years 1990–2007 (in-sample). The pricing accuracy of each estimated model is evaluated out-of-sample using the remaining 14 years of data, 2008–2021, in terms of the RMSE for log prices. We also present results for subsets of options, partitioning them by moneyness, days to maturity, or the level of VIX.

### v4

@@ 4.1 Data
Our empirical analysis is based on daily returns for the S&P 500 index, the CBOE VIX, and the panel of SPX option prices based on the S&P 500 index. Our sample period spans 32 years from January 2nd, 1990, to December 31, 2021, with 8,064 trading days. Returns are defined from cum-dividend logarithmic transformed closing prices, which were downloaded from the CRSP of Wharton Research Data Services (WRDS). Daily VIX prices were obtained from the Chicago Board Options Exchange (CBOE) website. The SPX option prices were obtained from two sources. Prices for the first six years (1990–1995) are the so-called Optsum data, which were purchased from CBOE website. Option prices for the remaining 26 years are OptionMetrics data downloaded from the WRDS database.
Option prices were primarily preprocessed following Christoffersen et al., (2014) and Bakshi et al., (1997). Specifically, we include out-of-the-money put and call option prices with positive trading volume and maturities between two weeks and six months. Options with missing implied volatilities or prices below one dollar are excluded. Put option prices are converted to call options using the put-call parity. Very deep out-of-the-money options with deltas larger than 0.85 or less than 0.15 are discarded. Much of the existing literature uses weekly data, and a panel of option prices is typically sampled on Wednesdays, because liquidity tends to be highest on Wednesdays. However, in our analysis, we use daily option prices, because a daily score, $s_{t}$, is needed to update $\eta_{t}$. The number of available options has grown rapidly over the sample period, both in terms of available maturities and the range of moneyness at each maturity. [^13] The number of available option prices has increased almost 50-fold over our sample period, initially from about 38 daily option prices to well over 1,700. From the pool of available options, we select up to six options per trading day. The inclusion criteria are as follows: we first determine the most liquid option for each maturity on day $t$, as measured by daily trading volume. We sort these options by maturity in ascending order and index these by $j=1,\ldots,\tilde{N}_{t}$, where $\tilde{N}_{t}$ is the number of distinct maturities on day $t$. From this set, we include all options if $\tilde{N}_{t}\leq 6$; the first six options if $7\leq\tilde{N}_{t}\leq 10$; options $\{1,3,5,7,9,11\}$ if $11\leq\tilde{N}_{t}\leq 15$; and options $\{1,4,7,10,13,16\}$ if $16\leq$ $\tilde{N}_{t}\leq 20$; and so forth. This resulting set of options will be representative for the range of available maturities, and we have $N_{t}=6$ on most days. The total number of option prices in our full sample period is 37,152.
@@
Note: Summary statistics for close-to-close S&P 500 index log returns, CBOE VIX, and SPX option prices from January 1990 to December 2021. We report the sample mean (Mean), standard deviation (Std), skewness (Skew), kurtosis (Kurt), number of observations (Obs), for returns and the VIX. Option prices are based on closing prices of out-of-the-money call and put options. We report the average Black-Scholes implied volatility (IV), average price, and the number of option prices for different ranges of Moneyness, days to maturity (DTM), and VIX levels. Data sources: S&P 500 returns from CRSP, VIX from CBOE’s website. Option prices from Optsum data (1990–1995) and OptionMetrics (1996–2021).
Table 2 contains descriptive statistics for our S&P 500 returns and the CBOE VIX in Panel A, and option prices in Panel B. As expected, the sample average of the VIX (19.48%) is larger than the standard deviation for annualized returns (18.11%). This difference reflects the (average) negative volatility risk premium. The S&P 500 returns exhibit slight negative skewness and a high degree of kurtosis, whereas the VIX has positive skewness and slightly lower kurtosis. We also report summary statistics for option prices partitioned by moneyness, maturity, and the contemporaneous level of VIX. Options with deltas below 0.5 are out-of-the-money call options, and options with deltas above 0.5 are based on out-of-the-money put options. Deep out-of-the-money put options (i.e., deltas greater than 0.7) are expensive relative to out-of-the-money call options, reflecting the well-known volatility smirk. The implied volatility has a relatively flat term structure with respect to time to maturity. Unsurprisingly, the implied volatility increases with the VIX level, as shown at the bottom of Table 2, where option prices are partitioned by the contemporaneous level of the VIX.
@@ 4.2 Parameter estimation
We estimate the model with both constant and time-varying variance risk aversion, CHNG and DHNG, respectively. [^14] The corresponding results for the Auxiliary structure are presented in the Appendix S.3. Both specifications are estimated using two types of derivative prices, VIX data and the panel of option prices. Parameters are estimated by maximizing the joint log-likelihood function for the full sample period from January 1990 to December 2021. Each column reports the parameter estimates for the specification listed in the first row of Table 3. The type of derivatives, VIX or option prices, used in the estimation is indicated with [VIX] and [Opt], respectively. Robust standard errors are give in parentheses below the estimates. [^15] Following Christoffersen et al., (2013), we impose $\omega$ = 0 in estimation when the non-negativity constraint, which ensures positive variances, is binding. We also report the implied persistence of volatility under both $\mathbb{P}$ and $\mathbb{Q}$. These are given by $\pi^{\mathbb{P}}=\beta+\alpha\gamma^{2}$ and $\pi^{\mathbb{Q}}=\mathbb{E}^{\mathbb{Q}}[\beta_{t}^{*}+\alpha_{t}^{*}\gamma_{t-1}^{*2}]$, respectively, where the latter simplifies to $\pi^{\mathbb{Q}}=\beta^{*}+\alpha^{*}\gamma^{*2}$ for the CHNG model with constant parameters. We also report the different terms of the log-likelihood (for returns and different sets of derivative prices). Some of these likelihood terms can be interpreted as pseudo out-of-sample log-likelihood values, indicated in italic. For instance, the CHNG[VIX] model is estimated using returns and the VIX, but the estimated model also yields model-based option prices that can be compared with actual option prices. The reported pseudo log-likelihood for option prices is evaluated with the resulting option pricing errors that are implicitly used to obtain estimates of $\rho$ and $\sigma_{e}$ for option prices. Similarly, we evaluate the log-likelihood for VIX pricing errors for the specifications estimated with option prices. In this case, we compute the implied estimate of $\sigma_{e}$ for VIX prices. The largest log-likelihood within each row is highlighted in bold.
@@
There are several interesting observations to be made from Table 3. First, the volatility process is found to be highly persistent across all specifications, and the persistence is larger under $\mathbb{Q}$ than under $\mathbb{P}$, which is consistent with the existing literature. Second, the estimate of the equity risk premium, $\lambda$, is positive and significant in all specification. Third, the VRR, $\eta$, is significantly larger than one in both CHNG specifications. Similarly, for the DHNG model the expected value of $\log\eta_{t}$, $\zeta$, is significantly positive in both DHNG specifications. Their corresponding expected $\eta_{t}$ (denoted $\mathbb{E}\eta$) inferred from the AR(1) model for $\log\eta_{t}$ are both notably larger than one. This implies that the risk-neutral volatility, $h^{\ast}$, is (on average) larger than the physical volatility, $h$. Fourth, the parameter that defines the leverage effects, $\gamma$, is also found to be significant across all specification. Fifth, for the DHNG specification we note that, $\eta_{t}$, is highly persistent, as $\varphi$ is estimated to be close to one. In fact, it is estimated to be more persistent than $h_{t}$ and $h_{t}^{\ast}$ across all DHNG specifications. Sixth, the coefficient, $\sigma$, which measures the impact of $s_{t-1}$ on $\eta_{t}$, is estimated to be positive and significant. [^16] In principle, $\sigma$ could be negative, but $\operatorname{var}(\sigma s_{t})=\sigma^{2}$ is non-negative. The implied unconditional variance for $\eta_{t}$ is similar for the two DHNG specifications, $\mathrm{var}(\eta_{t})=0.29$ when estimated with VIX and $\mathrm{var}(\eta_{t})=0.23$ when estimated with option prices. [^17] The variance for $\eta_{t}$ is computed from AR(1) parameters for $\log\eta_{t}$ and the assumptions imply that the unconditional distribution for $\eta_{t}$ is log-normal. Seventh, all specifications yield similar likelihood values for the returns. This is to be expected because they all rely on the same Heston-Nandi GARCH model for returns under physical measure.
@@
Eighth, the key difference between the CHNG and DHNG models is the enhanced flexibility in DHNG’s pricing kernel. This generalization leads to large improvements in the likelihood for derivative prices. The reason is that the adaptive pricing kernel with time-varying variance risk aversion yields model-implied derivative prices that are much closer to observed prices, and the substantial reduction in the pricing errors translates to much higher values of the log-likelihood for derivative prices. Ninth, estimating the models with VIX or option prices results in some differences across the estimated parameters. This is to be expected, as the vector of option prices contains more information about the distribution of future returns than the VIX. For instance, the leverage parameter, $\gamma$, is estimated to be larger with option prices than with the VIX. Tenth, the correlation among option pricing errors, $\rho$, is estimated to be positive and significant for both models. The estimate has the staggering large value of $82.2\%$ for the CHNG model and a more moderate value of $8.7\%$ for DHNG model. Ignoring this correlation (assuming it to be zero) would greatly underestimate the condition variance of $\nabla_{t}$, which makes the variance of $s_{t}$ larger than one.
@@
Figure 5: This figure displays the full-sample time series of the daily estimated variance risk ratio, $\eta_{t}$, from January 1990 to December 2021. The blue line is based on the VIX and the red line on option prices. The two time series are very similar, albeit the $\eta_{t}$ based on the VIX tend to be slightly larger than that based on option prices.
Alt text: Time series of the variance risk ratio from 1990 to 2021, estimated using VIX data and option prices. The two series follow similar trends, with occasional level differences, particularly during financial crises and market stress periods.
@@
Figure 5 presents the estimated daily time series of the VRR, $\eta_{t}=h_{t+1}^{\ast}/h_{t+1}$, based on the VIX (blue line) and option prices (red line). The two have a high degree of comovement and, as we discussed earlier, the unconditional variance of the two series is very similar. Importantly, the VRR occasionally falls below one, as is seen during the years 1993–1995, 2004–2007, and around 2017. A value below one ($\eta_{t}<1$) indicates that investors have an increased appetite for variance risk, whereas a large value of $\eta$ implies that investors demand additional compensation for variance risk. The latter was particularly the case during financial crises, such as the immediate aftermath of the Lehman collapse and the recent COVID-19 pandemic. The observed conditional variance was unusually high during these episodes, and the large value of $\eta_{t}$ implies that $h_{t}^{\ast}$ increased to far higher levels and was, briefly, more than three times larger than $h_{t}$.
@@
Note: This table presents the full-sample VIX and option pricing performance from January 1990 to December 2021 for each model listed in 3. We evaluate the model’s VIX and option pricing ability through the root mean square error for $\log{\rm VIX}$ and log implied volatility ($\log{\rm IV}$), respectively, see (10) and (11). Results for option pricing are presented for different ranges of moneyness (Black-Scholes Delta), days-to-maturity (DTM), and the VIX levels.
@@ 4.3 In-Sample VIX and Option Pricing Performance
We evaluate the performance of the models in terms of their ability to accurately price the VIX and options. We report the root mean squared errors (RMSE) for the logarithm of VIX and the logarithmically transformed implied volatilities. [^18] As a robustness check, we also computed the RMSE for the level of VIX and implied volatilities. These results are reported in Appendix S.1. The RMSE multiplied by 100 can be interpreted as the (absolute) relative pricing errors in percent. This loss function is coherent with our log-likelihood function, where the term that involves derivative prices is given in Assumption 4. A similar loss function was adopted in Ornthanalai, (2014), who used the relative implied volatility to prevent days with high volatility from being given a disproportionately large weight in the comparisons. For each of the specifications listed in Table 3, we report their (in-sample) RMSEs in Table 4.
@@
where $\mathrm{VIX}_{t}^{m}$ is the model-based quantity and $\mathrm{VIX}_{t}$ is the observed market VIX value. Panel A of Table 4 shows that a time-varying VRR, $\eta_{t}$, results in substantially smaller average pricing errors, relative to a constant $\eta$, which is a characteristic of CHNG. The improved VIX pricing is substantial and impressive. The most accurate VIX pricing is achieved with by the DHNG model, when it is estimated with VIX data. This reduces the RMSE by a factor of four relative to both CHNG specifications. Even the DHNG model that is estimated solely with option prices, has a much smaller RMSE for VIX prices than the CHNG specification that uses VIX prices as part of the objective in the estimation. This is impressive, because the DHNG model estimated with option prices is based on an objective that simultaneously seeks to price options with maturities ranging from one to six months. This entails a trade-off across maturities, unlike models estimated with solely with VIX, that specifically targets the one-month maturity.
@@
Figure 6: This figure displays the full-sample time series of the daily model-based (red lines) and market-based (blue lines) VIX and average implied volatility. The CHNG and DHNG models are estimated with VIX data in the upper panels and option prices in the lower panels. Model parameters are based on the Table 3.
Alt text: Comparison of model-based and market-based values for the VIX and average implied volatility over time. The figure presents results for both the CHNG and DHNG models, estimated using VIX data in the upper panels and option prices in the lower panels. DHNG provides a closer match to market values.
@@
Note: Each model is estimated once, using data from the years 1990–2007 (in-sample). The pricing accuracy of each estimated model is evaluated out-of-sample using the remaining 14 years of data, 2008–2021, in terms of the RMSE for log prices. We also present results for subsets of options, partitioning them by moneyness (Black-Scholes Delta), days-to-maturity (DTM), and the VIX levels.